OpenStudy (princezz):

F(x)=x^2-4/x+2 What is the domain and the vertical asymptote?

OpenStudy (anonymous):

the domain is everything except where \(x+2=0\) i.e at \(x=-2\)

OpenStudy (princezz):

Oh ok that what I got too as my answer

OpenStudy (anonymous):

there are no vertical asymptotes, factor and cancel and you see \[\frac{x^2-4}{x+2}=\frac{(x+2)(x-2)}{x+2}=x-2\]